# On the Hardness of Approximation of the Fair $k$-Center Problem

### v1

@@ Abstract
In this work, we study the hardness of approximation of the fair $k$-center problem. Here the data points are partitioned into groups and the task is to choose a prescribed number of data points from each group, called centers, while minimizing the maximum distance from any point to its closest center. Although a polynomial-time $3$-approximation is known for this problem in general metrics, it has remained open whether this approximation guarantee is tight or could be further improved, especially since the unconstrained $k$-center problem admits a polynomial-time factor-$2$ approximation. We resolve this open question by proving that, for every $\epsilon>0$, achieving a $(3-\epsilon)$-approximation is $\mathsf{NP}$-hard, assuming $\mathsf{P}\neq\mathsf{NP}$.
@@ 1 Introduction
Clustering is a fundamental problem in data analysis and optimization, with applications ranging from uncovering structure in large datasets to selecting representatives for data summarization, and supporting facility-location decisions [23, 3]. In its standard form, clustering involves selecting $k$ representative points—called centers—and assigning every data point to a chosen center, thereby inducing a partition of the data into $k$ clusters. The quality of the resulting solution is measured by an objective that quantifies how well the centers “cover” the data. A particularly well-studied objective is the $k$-center problem, which seeks $k$ centers that minimize the maximum distance from any point to its nearest center. This min-max distance criterion captures a natural “worst-case” notion of quality—closely related to an egalitarian fairness principle, since it focuses on protecting the most poorly served point. Though the problem is $\mathsf{NP}$-hard, it admits a simple and practical constant-factor approximation algorithm: starting with a random point as a center and greedily choosing farthest point from the chosen centers (for k1k-1 iterations) achieves a $2$-approximation and this guarantee is essentially tight, assuming $\mathsf{P}\neq\mathsf{NP}$ [13, 16]. Owing to its practicality, $k$-center has become a canonical testbed for understanding both the algorithmic possibilities and the inherent limitations of center-based clustering.
@@
When data points are associated with sensitive attributes that induce demographic groups, and clustering is used to select representatives—for example, when the chosen centers are presented as a summary of the data, or used to serve as exemplars for downstream decision-making—optimizing a standard clustering objective alone can yield biased representation for certain groups [23]. A growing body of work has therefore investigated fair clustering that explicitly encode desiderata related to sensitive attributes via explicit representation constraints [30, 28, 29, 10, 23, 21, 15, 24]. A particularly natural notion for data summarization is representative (or center-based) fairness, which requires the set of chosen centers to include a prescribed numbers of representatives from each group, termed as the fair $k$-center problem [23, 21]. We emphasize that fair data summarization is studied in two closely related formulations: $k$-center and $k$-supplier. In the $k$-center setting, any data point can be chosen as a center and every point is also a client that must be served. In the $k$-supplier setting, centers must be chosen from a designated subset of points, called facilities, which may differ from the client set. This distinction is subtle but consequential: in $k$-supplier, facilities are not required to be served by the opened centers, whereas in $k$-center every data point, including potential center locations, must be served [10, 4].
@@
On the algorithmic side, fair $k$-center admits constant-factor approximations in general metrics. The current best-known approximation guarantee is factor-$3$: first compute a “fairness-oblivious” set of candidate centers using Gonzalez’s greedy procedure achieving factor-$2$ approximation, and then project these candidates to a set of “fair centers” that satisfy the group requirements via a matching-based step. The analysis bounds the extra cost incurred during this projection, yielding an overall $3$-approximation [21]. Strikingly, the picture has remained incomplete. While a factor-$3$ approximation is known for fair $k$-center, it was not known whether this guarantee is tight. This uncertainty is particularly salient when contrasted with the classical (unconstrained) $k$-center problem, which admits a polynomial-time $2$-approximation and for which this factor-$2$ is essentially optimal. It is therefore natural to ask whether fairness constraints inherently force a strictly worse approximation factor, or whether we could hope to “recover” the factor-$2$ guarantee in the fair setting. The comparison to the $k$-supplier variant further sharpens the question: both $k$-supplier and fair $k$-supplier admit polynomial-time $3$-approximation, and this factor is tight in both cases, so fairness does not create an approximation gap there. For $k$-center, by contrast, an apparent gap persists between the $2$-approximation for the unconstrained variant and the $3$-approximation for the fair variant. The main goal of this work is to determine whether this gap can be bridged, i.e., whether the problem has a genuine computational barrier or this is an artifact of current algorithmic techniques employed to solve this problem.
@@
Understanding these limitations are important not only from a theoretical stand point but also in practice. Hardness of approximation results clarify the algorithmic limits of optimization problems such as (fair) clustering: they delineate which improvements are achievable and help avoid chasing approximation guarantees that are provably unattainable. Equally importantly, they sharpen the research agenda by indicating where progress may still be possible—for example, by exploiting additional structure in metric spaces, e.g., Euclidean or doubling metrics, or by considering meaningful relaxations to fairness constraints. In this paper, we extend this line of research, by investigating the hardness of approximation of the fair $k$-center problem.
@@
Technical challenges. Our hardness of approximation results build upon the inapproximability of the $k$-center problem with costs. Specifically, Chuzhoy et al. [7][Theorem 4.1] show that $k$-center with costs is $\mathsf{NP}$-hard to approximate within a factor of $3-\epsilon$, for any $\epsilon>0$. This hardness carries over to the closely related $k$-center problem with forbidden centers [1], in which the set of data points $U$ is partitioned into two groups $A\cup B$ and it is allowed to open centers only in $A$, in other words, points in $B$ are forbidden as center locations. [^1] Note that, in their $(3-\epsilon)$-inapproximability proof for $k$-center with costs, Chuzhoy et al. [7][Theorem 4.1] assign unit cost to the “regular” points and prohibitively large costs to a designated subset of points. This choice effectively rules out those points with larger costs as possible centers while still requiring them to be covered by a opened center. Consequently, their construction can be interpreted as an instance of the $k$-center with forbidden centers: some points must be served, but are not allowed to be selected as centers. This formulation is different from $k$-supplier, where the facility set is distinct from the client set, and facilities themselves need not be served by any opened center. In contrast, in $k$-center with forbidden centers (equivalently, a ”$k$-supplier“ view with forbidden locations), points in $B$ are disallowed as center locations, but they remain part of the metric and therefore must still be assigned to some opened center in $A$.
@@
Observe that $k$-center with forbidden centers is a special case of fair $k$-center with two groups $G_{1}=A,G_{2}=B$, and requirements $r_{1}=k$, $r_{2}=0$. This implies a immediate $(3-\epsilon)$-inapproximability in the degenerate regime where some groups can have zero requirement. In contrast, many fair clustering settings assume that the data is partitioned into disjoint (demographic) groups $\mathbb{G}={G_{1},\dots,G_{t}}$ with $G_{i}\cap G_{j}=\emptyset$ and $\bigcup_{i\in[t]}G_{i}=U$; for instance, in the two-group setting one may partition points by sex (i.e., Male and Female). In such scenarios, allowing $r_{i}=0$ effectively permits selecting no centers from an entire demographic group—an outcome that is usually inconsistent with the intended fairness goal and arguably too weak to serve as evidence of inapproximability for a practically relevant problem. Most importantly, the above reduction does not establish inapproximability in the regime where every group must be represented among the centers, i.e., when the requirements are $1\leq r_{i}\leq k-1$.
@@
Moreover, several prior works [30, 29, 10, 28, 11, 8, 9] use a common reduction paradigm: they transform fair clustering instances with arbitrary group requirements into a simpler canonical form consisting of exactly $k$ disjoint groups, with the constraint that the solution must select exactly one center from each group. Many fairness constraints and clustering variants admit such a reduction, so an efficient algorithm for this “one-per-group” setting would immediately yield efficient algorithms for the more general formulations. This approach has been particularly powerful for objectives such as $k$-median, $k$-means, and $k$-supplier: after reducing to the one-per-group case, one can obtain approximation guarantees that match those of the corresponding unconstrained problems.
@@
This raises a natural question: is fair $k$-center $\mathsf{NP}$-hard to approximate within a factor strictly better than $3$ in polynomial time when the fairness constraints are non-degenerate—specifically, when exactly one center must be chosen from each group?
@@
Our contributions. We establish hardness of approximation for fair $k$-center, resolving an open problem by showing that the factor-$3$ approximation achievable in polynomial time is essentially optimal. Our results cover two non-degenerative cases: ($i$) instances with two groups, where at least one center must be selected from each group, and ($ii$) instances with an arbitrary number of groups in the one-per-group regime, where exactly one center must be selected from each group.
@@
The remainder of the paper is organized as follows. In Section 2 we introduce the problem definitions. Section 3 presents our hardness-of-approximation results for fair $k$-center. Section 4 reviews the most relevant related work, and Section 5 concludes with a discussion of implications and possible directions of future work.
@@ 2 Problem definitions
$\text{minimize}\quad\texttt{cost}(S):=\max_{u\in U}\min_{s\in S}d(u,s).$
@@
Definition 3 (The $k$-center problem with forbidden centers).
@@
$\text{minimize}\quad\texttt{cost}(S):=\max_{u\in U}\;\min_{s\in S}d(u,s).$
@@ 3 Hardness of approximation
In this section, we present our hardness-of-approximation results. We first show that fair $k$-center with two groups remains $\mathsf{NP}$-hard to approximate within a factor of $3-\epsilon$, for any $\epsilon>0$, even when the requirements are non-degenerate, i.e., when at least one center must be selected from each group. We then extend this hardness to the canonical one-per-group regime by transforming any two-group fair $k$-center instance into an equivalent instance with $k$ groups with a requirement of choosing one-center per group. The transformation duplicates points to create the desired number of groups, and we preserve the metric structure by introducing a small separation $\delta>0$ between duplicates. By choosing $\delta$ appropriately, we ensure that, for every $\epsilon>0$, a $(3-\epsilon)$-approximation for the one-per-group variant would imply a $(3-\epsilon)$-approximation for the two-group case, and hence cannot exist unless $\mathsf{P}=\mathsf{NP}$.
@@ 3.1 Two groups regime
We prove the claimed inapproximability result via a polynomial-time reduction from $k$-center with forbidden centers to fair $k$-center with two groups. Our argument proceeds in three steps. ($i$) We construct, in polynomial time, a fair $k$-center instance with two groups by adding a single auxiliary point $x$ and defining its distances to all other points so that $x$ is effectively isolated (1). ($ii$) We prove that every optimal solution to the resulting fair instance must include $x$ as a center (2), and that removing $x$ preserves the objective value on the remaining points, establishing an equivalence between optimal solutions of the two problems (3). (iiiiii) We show that any $\alpha$-approximate solution for the fair instance can be converted into an $\alpha$-approximate solution for the original forbidden-centers instance by simply discarding $x$ (4). Combining these steps with the known $(3-\epsilon)$-inapproximability of $k$-center with forbidden centers yields our hardness of approximation for fair $k$-center.
@@
Note that the inapproximability result of Theorem 1 extends to fair $k$-center with two groups with an arbitrary (non-zero) requirement for each group. The only modification in the construction of 1 is to introduce $r_{2}\geq 1$ auxiliary points $\{x_{1},\dots,x_{r_{2}}\}$ and place them in group $G_{2}$ (i.e., $G_{2}=B\cup\{x_{1},\dots,x_{r_{2}}\}$), and $U^{\prime}=U\cup\{x_{1},\dots,x_{r_{2}}\}$. Furthermore, set distances in $d^{\prime}$ as follows:
$d(u,v)$ $=0,\quad\text{for all}\quad u,v\in U,$
@@
$d^{\prime}(x_{i},u)=d^{\prime}(u,x_{i})$ $=3D+1,\quad\text{for each}\quad x_{i}\in\{x_{1},\dots,x_{r_{2}}\},\;u\in U^{\prime}\setminus\{x_{i}\}$

### v2

@@ Abstract
In this work, we study the hardness of approximation of the fair $k$-center problem. In this problem, we are given a set of data points in a metric space that is partitioned into groups and the task is to choose a subset of $k$-data points, called centers, such that a prescribed number of data points from each group are chosen while minimizing the maximum distance from any point to its closest center. Although a polynomial-time $3$-approximation is known for fair $k$-center in general metrics, it has remained open whether this approximation guarantee is tight or could be further improved, especially since the classical unconstrained $k$-center problem admits a polynomial-time factor-$2$ approximation. We resolve this open question by proving that, assuming $\mathsf{P}\neq\mathsf{NP}$, for any $\epsilon>0$, no polynomial-time algorithm can approximate fair $k$-center to $(3-\epsilon)$-factor.
@@ 1 Introduction
Clustering is a fundamental problem in data analysis and optimization, with applications ranging from uncovering structure in large datasets to selecting representatives for data summarization, and supporting facility-location decisions [24, 4, 28, 16, 17, 22]. In its standard form, clustering involves selecting $k$ representative points—called centers—and assigning every data point to a chosen center, thereby inducing a partition of the data into $k$ clusters. The quality of the resulting solution is measured by an objective that quantifies how well the centers “cover” the data points. A particularly well-studied objective is the $k$-center problem, which seeks $k$ centers that minimize the maximum distance from any point to its nearest center. This min-max distance criterion captures a natural “worst-case” notion of quality—closely related to an egalitarian fairness principle, since it focuses on protecting the most poorly served point. Though the problem is $\mathsf{NP}$-hard, it admits a simple and practical constant-factor approximation algorithm: starting with a random point as a center and greedily choosing farthest point from the chosen centers (for k1k-1 iterations) achieves a $2$-approximation and this guarantee is essentially tight, assuming $\mathsf{P}\neq\mathsf{NP}$ [14, 17]. Owing to its practicality, $k$-center has become a canonical testbed for understanding both the algorithmic possibilities and the inherent limitations of center-based clustering.
@@
When data points are associated with sensitive attributes that induce demographic groups, and clustering is used to select representatives—for example, when the chosen centers are presented as a summary of the data, or used to serve as exemplars for downstream decision-making—optimizing a standard clustering objective alone can yield biased representation for certain groups [3, 6]. A growing body of work has therefore investigated fair clustering that explicitly encode desiderata related to sensitive attributes via explicit representation constraints [31, 29, 30, 11, 24, 22, 16, 25]. A particularly natural notion for data summarization is representative (or center-based) fairness, which requires the set of chosen centers to include a prescribed numbers of representatives from each group, termed as the fair $k$-center problem [24, 22]. We emphasize that (fair) data summarization is studied in two closely related formulations: $k$-center and $k$-supplier objectives. In the (fair) $k$-center setting, any data point can be chosen as a center and every point is also a client that must be served. In the (fair) $k$-supplier setting, centers must be chosen from a designated subset of points, called suppliers, which may differ from the client set. This distinction is subtle but consequential: in (fair) $k$-supplier, suppliers are not required to be served by the opened centers, whereas in (fair) $k$-center every data point, including potential center locations, must be served [11, 5, 14, 18].
@@
On the algorithmic side, fair $k$-center admits constant-factor approximation algorithms in general metrics. The current best-known approximation guarantee is factor-$3$: first compute a “fairness-oblivious” set of candidate centers using Gonzalez’s greedy procedure achieving factor-$2$ approximation, and then project these candidates to a set of “fair centers” that satisfy the group requirements via a matching-based step. The analysis bounds the extra cost incurred during this projection, yielding an overall $3$-approximation [22]. Strikingly, the picture has remained incomplete. While a factor-$3$ approximation is known for fair $k$-center, it was not known whether this guarantee is tight. This uncertainty is particularly salient when contrasted with the classical (unconstrained) $k$-center problem, which admits a polynomial-time $2$-approximation and for which this factor-$2$ is essentially optimal. It is therefore natural to ask whether fairness constraints inherently force a strictly worse approximation factor, or whether we could hope to “recover” the factor-$2$ guarantee in the fair setting. The comparison to the $k$-supplier variant further sharpens the question: both unconstrained $k$-supplier and fair $k$-supplier admit polynomial-time $3$-approximation, and this factor is tight in both cases, so fairness does not create an approximation gap there. In contrast, for $k$-center an apparent gap persists between the $2$-approximation for the unconstrained variant and the $3$-approximation for the fair variant. The main goal of this work is to determine whether this gap can be bridged, i.e., whether the problem has a genuine computational barrier or this is an artifact of current algorithmic techniques employed to solve this problem.
@@
Understanding these limitations are important not only from a theoretical stand point but also in practice. Hardness of approximation results clarify the algorithmic limits of optimization problems such as (fair) clustering: they delineate which improvements are achievable and help avoid chasing approximation guarantees that are provably unattainable. Equally important, they sharpen the research agenda by indicating where progress may still be possible—for example, by exploiting additional structure in metric spaces, e.g., Euclidean or doubling metrics, or by considering meaningful relaxations to fairness constraints. In this paper, we extend this line of research, by investigating the hardness of approximation of the fair $k$-center problem.
@@
Technical challenges. Our hardness of approximation results build upon the inapproximability of the $k$-center problem with costs. Specifically, Chuzhoy et al. [8][Theorem 4.1] show that $k$-center with costs is $\mathsf{NP}$-hard to approximate within a factor of $3-\epsilon$, for any $\epsilon>0$. In this variant, each point has a nonnegative opening cost, and the total cost of chosen centers should not exceed a given budget. The hardness of $k$-center with costs carries over to the closely related $k$-center with forbidden centers problem [1], in which the set of data points $U$ is partitioned into two groups $A\cup B$ and it is allowed to open centers only in $A$, in other words, points in $B$ are forbidden as center locations. [^1] Note that, in their $(3-\epsilon)$-inapproximability proof for $k$-center with costs, Chuzhoy et al. [8][Theorem 4.1] assign unit cost to the “regular” points and prohibitively large costs (larger than $k$ suffices) to a designated subset of points. This choice of cost assignment effectively rules out those points with larger costs as possible centers while still requiring them to be covered by an opened center. Consequently, their construction can be interpreted as an instance of $k$-center with forbidden centers: some points must be served, but are not allowed to be selected as centers. Angelidakis et al. [1] makes this connection explicit. Citing Chuzhoy et al. [8], they note that $k$-center with forbidden centers is $\mathsf{NP}$-hard to approximate to a factor $(3-\epsilon)$, and argue that $k$-center with forbidden centers is a special case of the matroid center problem. Consequently, the same $(3-\epsilon)$ inapproximability transfers to matroid center as well. Although their discussion is framed around hardness of approximation for matroid center, the underlying source of the barrier is precisely the $(3-\epsilon)$-inapproximability of $k$-center with forbidden centers (or costs) from Chuzhoy et al. [8][Theorem 4.1]. This formulation is different from $k$-supplier, where the facility set can be distinct from the client set, and suppliers themselves need not be served by any opened center. In contrast, in $k$-center with forbidden centers (equivalently, a “$k$-supplier” view with forbidden locations), points in $B$ are disallowed as center locations, but they remain part of the metric and therefore must still be assigned to some opened center in $A$.
@@
Observe that $k$-center with forbidden centers is a special case of fair $k$-center, with two groups $G_{1}=A$, $G_{2}=B$ and requirements $r_{1}=k$, $r_{2}=0$. This implies an immediate $(3-\epsilon)$-inapproximability in the degenerate regime where some groups can have zero requirement. In contrast, many fair clustering settings assume that the data is partitioned into disjoint (demographic) groups $\mathbb{G}=\{G_{1},\dots,G_{t}\}$ with $G_{i}\cap G_{j}=\emptyset$ and $\bigcup_{i\in[t]}G_{i}=U$; for instance, in the two-group setting one may partition points by sex (i.e., Male and Female). In such scenarios, allowing $r_{i}=0$ effectively permits selecting no centers from an entire demographic group—an outcome that is usually inconsistent with the intended fairness goal and arguably too weak to justify inapproximability claims for a practically relevant problem, which extends beyond such degenerate cases. Most importantly, the above reduction does not establish inapproximability in the regime where every group must be represented among the centers, i.e., when the requirements are $1\leq r_{i}\leq k-1$.
@@
Moreover, several prior works [31, 30, 11, 29, 12, 9, 10] use a common reduction paradigm: they transform fair clustering instances with arbitrary group requirements into a simpler canonical form consisting of exactly $k$ disjoint groups, with the constraint that the solution must select exactly one center from each group. Many fairness constraints and clustering variants admit such a reduction, so an efficient algorithm for this “one-per-group” setting would immediately yield efficient algorithms for the more general formulations. This approach has been particularly powerful for objectives such as $k$-median, $k$-means, and $k$-supplier: after reducing to the one-per-group case, one can obtain approximation guarantees that match those of the corresponding unconstrained problems. This raises a natural question.
@@
Question: Is fair $k$-center $\mathsf{NP}$-hard to approximate within a factor strictly better than $3$ in polynomial time when the fairness constraints are non-degenerate—specifically, when exactly one center must be chosen from each group?
@@
Our contributions. We establish hardness of approximation for fair $k$-center and resolve an open problem, by showing that the factor-$3$ approximation achievable in polynomial time is essentially optimal. Our results cover two non-degenerative cases: ($i$) instances with two groups, where at least one center must be selected from each group, and ($ii$) instances with an arbitrary number of groups in the one-per-group regime, where exactly one center must be selected from each group.
@@
The remainder of the paper is organized as follows. In Section 2 we introduce the problem definitions. Section 3 presents our hardness of approximation results. Section 4 reviews the most relevant related work, and Section 5 concludes with a discussion of implications of our inapproximability results and possible directions of future work.
@@ 2 Problem Definitions
$\text{minimize}\quad\texttt{cost}(S,U):=\max_{u\in U}\min_{s\in S}d(u,s).$
@@
Definition 3 (The $k$-center with forbidden centers problem).
@@
$\text{minimize}\quad\texttt{cost}(S,U):=\max_{u\in U}\;\min_{s\in S}d(u,s).$
@@ 3 Hardness of Approximation
In this section, we present our hardness of approximation results. We first show that fair $k$-center with two groups remains $\mathsf{NP}$-hard to approximate within a factor of $3-\epsilon$, for any $\epsilon>0$, even when the requirements are non-degenerate, i.e., when at least one center must be selected from each group. We then extend this hardness to the canonical one-per-group regime by transforming any fair $k$-center with two groups instance into an equivalent instance with $k$ groups with a requirement of choosing exactly one-center per group. The transformation duplicates points to create the desired number of groups, and we preserve the metric structure by introducing a small separation $\delta>0$ between duplicates. By choosing $\delta$ appropriately, we ensure that, for every $\epsilon>0$, a $(3-\epsilon)$-approximation for the one-per-group variant would imply a $(3-\epsilon)$-approximation for the two-group case, and hence such an approximation algorithm cannot exist.
@@ 3.1 Two groups regime
We prove the claimed inapproximability result via a polynomial-time reduction from $k$-center with forbidden centers to fair $k$-center with two groups. Our argument proceeds in three steps. ($i$) We construct, in polynomial time, a fair $k$-center instance with two groups by adding a single auxiliary point $x$ and defining its distances to all other points so that $x$ is effectively isolated (1). ($ii$) We prove that every optimal solution to the resulting fair instance must include $x$ as a center (2), and that removing $x$ preserves the objective value on the remaining points, establishing an equivalence between optimal solutions of the two problems (3). (iiiiii) We show that any $\alpha$-approximate solution for the fair $k$-center with two groups instance can be converted into an $\alpha$-approximate solution for $k$-center with forbidden centers instance by simply discarding $x$ (4). Combining these steps with the known $(3-\epsilon)$-inapproximability of $k$-center with forbidden centers yields our hardness of approximation for fair $k$-center with two groups.
@@
Note that the inapproximability result of Theorem 1 extends to fair $k$-center with two groups with an arbitrary (non-zero) requirement for each group. The only modification in the construction of 1 is to introduce $r_{2}\in\{1,\dots,k-1\}$ auxiliary points $\{x_{1},\dots,x_{r_{2}}\}$ and place them in group $G_{2}$ (i.e., $G_{2}=B\cup\{x_{1},\dots,x_{r_{2}}\}$), $U^{\prime}=U\cup\{x_{1},\dots,x_{r_{2}}\}$, and then set $r_{1}=k$, and $k^{\prime}=k+r_{2}$. Furthermore, set distances in $d^{\prime}$ as follows:
@@
$d^{\prime}(u,v)$ $=d(u,v),\quad\text{for all}\quad u,v\in U,$
$d^{\prime}(x_{i},u)=d^{\prime}(u,x_{i})$ $=3D+1,\quad\text{for each}\quad x_{i}\in\{x_{1},\dots,x_{r_{2}}\},\;\text{and}\quad u\in U^{\prime}\setminus\{x_{i}\}.$
@@
Remark 2.
We present our hardness of approximation result in the two-group regime to highlight that the $(3-\epsilon)$-inapproximability already holds when the number of groups is small. Moreover, such two-group instances can be transformed into equivalent instances with more groups (up to $k$) by duplicating points and adjusting the per-group requirements accordingly. Hence, the same $(3-\epsilon)$-inapproximability extends to settings with an arbitrary number of groups; in this sense, the two-group regime is a strictly restrictive case. This transformation using duplication step is exactly what we use to establish the hardness result for the one-per-group fair $k$-center problem, which we explain in the subsection to follow.